# Beyond N-gram: Data-Aware X-gram Extraction for Efficient Embedding Parameter Scaling

### 23 Apr 2026 v1

@@ 1 Introduction
Increasing model capacity in large language models usually comes with higher inference-time compute, making further scaling increasingly costly in practice. Token-indexed memory offers an alternative scaling path: instead of relying solely on dense computation, the model can retrieve token-conditioned features from a lookup table, thereby expanding static capacity without proportionally increasing compute [1, 2]. In practice, however, prior lookup-based designs often require very large tables and substantial memory traffic, making these gains difficult to realize efficiently [1, 2].
@@
To make this scaling axis practical, we ask a more basic question: what does an efficient lookup table need to optimize for under finite data and bandwidth? We analyze how token-indexed parameters are trained and how their representations behave, and find two consistent regularities. First, utilization is strongly frequency-driven under Zipf-like statistics: high-frequency items dominate updates and learned activations, while most long-tail rows remain under-trained as the table grows [3]. Second, scaling capacity by adding parallel fixed slots is not sufficient: different slots are trained on highly overlapping events, encouraging redundant subspaces and rapid saturation [1]. These findings indicate that the bottleneck is not the lack of parameters, but the mismatch between table capacity, update density, and representational diversity.
@@ 2 General Framework
Lookup-based augmentation is fundamentally an efficiency problem: the injection design determines how much memory, activation traffic, and compute are spent per token. We therefore track three budgets: (i) memory budget (cache size) $\mathcal{C}$, dominated by the lookup table size $S\cdot d_{s}$ where $S$ is the number of physical rows and $d_{s}$ is the injected dimensionality at injection site $s$; (ii) activation budget $\mathcal{A}$, the per-token activation footprint of the injected pathway, proportional to $d_{s}$; Here, activation refers to the memory footprint of intermediate tensors saved for backpropagation, rather than nonlinear activation functions. and (iii) compute budget $\mathcal{F}$, the extra per-token FLOPs induced by lookup aggregation, feature extraction, and optional projections, which also scales with $d_{s}$. We refer to improvement per unit $\mathcal{C}$ as parameter efficiency, and improvement per unit $\mathcal{A}$ as activation efficiency.
@@ 3 Observation
Unless stated otherwise, we report statistics after training for 41.4B tokens; metric definitions are given in Appendix C.4.
@@
We then ask why simply adding more parallel lookup slots often yields diminishing returns even when memory budget increases. In fixed-slot designs, each slot sees almost the same token occurrences and thus receives highly correlated gradients; as a result, “more slots” does not reliably create “more directions” in representation space. We diagnose this by measuring pairwise cosine similarity between retrieved embeddings across slots (Appendix C.4). Figure 2 shows that slot representations are highly similar, indicating subspace collapse and limited angular diversity. The implication is that parameter count alone is not a sufficient scaling axis: without a mechanism that breaks symmetry across slots, additional tables mostly replicate existing features.
@@ 4.2 X-gram Extraction with Gated ShortConv
Lookup retrieval is context-independent, but whether a retrieved vector is useful depends on local context. Moreover, without a view-specific nonlinear extractor, parallel memories can collapse to redundant subspaces (Appendix C.6). We therefore refine each retrieved sequence with a lightweight, normalized, gated multi-scale ShortConv that extracts local $x$-gram cues and promotes diversity across views.
@@
Different views use different kernel sizes, so the resulting extractors collectively provide multi-scale coverage and expand the receptive field from token identity to local $x$-grams. The gate acts as a context-selective filter that suppresses collision-induced noise. View-specific kernels break symmetry across parallel memories, increasing diversity before fusion. Additional rationale is provided in Appendix C.7.
@@ 4.4 Training Stability and Optimization
Token-indexed parameters are updated sparsely and can destabilize training if their scale and learning dynamics are mismatched with the backbone. We control stability with three scale rules and a sparse-aware optimization scheme; detailed derivations are provided in Appendix C.5.
@@
For optimization, we treat lookup-table parameters as a sparse group and apply a larger effective step size than the backbone to compensate for low update frequency. Concretely, we scale the lookup learning rate as a function of expected row hit rate, which depends on table size $S$ (equivalently $\rho|\mathcal{V}|$); a practical rule is $\eta_{\text{lookup}}\propto\sqrt{S}$ up to clipping, with further refinements in Appendix C.5.
@@ 5.1 Experimental Setup
We compare against a vanilla decoder-only Transformer and representative compute-decoupled baselines that scale capacity via token-indexed parameters or embedding modules (e.g., STEM, MoLE, and SCONE-style conditional memory) [4, 1, 2]. All methods share the same backbone at each scale (0.73B and 1.15B) and the same training token budget; differences are restricted to the token-indexed parameterization, feature extraction, and injection design. Backbone hyperparameters and training recipes are summarized in Table 9 and Appendix B.2 (extended setup in Appendix B.1).
@@
We pretrain on OLMo-mix-1124 [5] (web + domain data from DCLM/Dolma) [6, 7] and evaluate with lm-evaluation-harness [8] on a fixed downstream suite with consistent prompts and shot counts (Table 12).
@@ 5.5 Related Work
A recent thread explores adding large token- or $n$-gram-indexed parameters as an auxiliary memory channel, aiming to scale capacity with minimal additional FLOPs. Mixture-of-lookup designs route tokens to a small set of retrieved vectors and combine them with lightweight mixing, enabling substantial parameter growth while keeping the runtime compute largely unchanged [1]. Similarly, scalable lookup-based conditional memory frames retrieval as a new axis of sparsity: additional capacity primarily increases memory traffic rather than matrix multiplication [2]. These methods are attractive in practice because retrieval is deterministic from token IDs and can be served from host memory with prefetching, but they also expose a core bottleneck: the efficiency of using those parameters per accessed byte, especially under Zipfian long tails where many rows are rarely updated.
@@
Another line of work enlarges the input interface by scaling embedding layers, vocabulary size, or related embedding modules [3, 9, 4]. Closely related work also studies memory-efficient embedding parameterizations through heterogeneous capacity allocation, compositional compression, and parameter sharing, especially in recommendation systems and large-vocabulary language models [10, 11, 12, 13, 14, 15, 16, 17]. These approaches mainly improve the capacity or memory efficiency of input-side embeddings, often through non-uniform allocation or shared parameterization. In contrast, we study layer-wise token-indexed memory rather than a single input embedding layer, and use frequency information to structure physical capacity allocation, with hashing applied only locally within dense buckets where collisions are unavoidable.
@@
Sparse Mixture-of-Experts (MoE) [18] employs learnable gating to activate a sparse subset of experts per token, enabling efficient capacity scaling [19, 20, 21]. While recent advances explore fine-grained specialization [22] and unified scaling laws [23], standard MoE designs typically partition experts by layer. This layer-wise isolation causes total memory to scale linearly with depth and limits the potential for cross-layer operator reuse. Existing routing strategies—ranging from load-balancing heuristics [20, 19] and hash-based assignments [24] to expert-choice [25], BASE layers [26], and soft mixtures [27]—primarily optimize intra-layer token allocation. However, these methods do not address the recursive, multi-layer exposure of experts, which is central to the universal expert design in X-gram.
@@
Where injection occurs matters for both stability and efficiency. Injecting into the attention value pathway can be particularly cost-effective under grouped-query attention because the value dimensionality is smaller than the query/key dimensionality, yielding lower activation and compute overhead for the same memory budget. Recent analyses of value-residual designs further highlight the importance of value-stream information flow [28]. Our method adopts value-path injection as the default site and uses depth-aware gating and normalization to keep the injected signal stable during training.
@@ 6 Conclusion
Appendix A Notation
@@
Appendix B Full Versions of Compressed Sections
B.1 Extended Experimental Setup
We provide additional experimental details that are compressed in Section 5.1. Baselines. We compare against (i) Vanilla Transformer, the standard decoder-only backbone, and representative compute-decoupled baselines that scale capacity via embedding modules or token-indexed memory, including STEM [4], MoLE [1], and SCONE [2]. All models share identical backbones (same depth/width) and are trained under the same token budget; methods differ only in the token-indexed parameterization and injection design. Data and evaluation. Models are pretrained on OLMo-mix-1124 [5], which integrates high-quality web text from DCLM [6] with diverse domain-specific data from Dolma [7]. We evaluate using lm-evaluation-harness [8] and keep prompts/shot counts fixed across methods within each benchmark (Table 12). Budget alignment. We keep additional compute from lookup and ShortConv within a small constant factor (Table 2) and align the memory budget $\mathcal{C}$ and injected views/slots across token-indexed methods (Table 11).
B.2 Training Details
@@
B.3 Observation (extended)
We conduct all observations by training two models under identical data and compute budgets: a Base Transformer and a PLE variant that injects token-indexed memory into every layer. In the PLE model, each layer aggregates $M_{\ell}$ token-indexed views and injects their normalized sum into the layer according to the PLE pipeline described in Section 4. Concretely, we use Eq. (5), and we report statistics after training for 41B tokens. We provide exact definitions for the activation norm and similarity metrics in Appendix C.4.
Trainability is dominated by Zipfian frequency. To sharpen the optimization target, we first ask a basic but overlooked question: when a token-indexed lookup table is scaled up, are its parameters actually trainable under realistic data budgets? In practice, the answer is constrained by Zipf’s law. Because token and $n$-gram sequences follow an extreme long tail, updates to large tables become highly sparse: most entries are rarely, if ever, visited, and therefore remain effectively under-trained. This under-training is not a minor edge case but a structural property of frequency distributions. For example, after processing 100M tokens, only $7.3\%$ of entries in a 2M-entry table receive more than 100 updates, whereas a 32K table reaches $97.6\%$ under the same criterion. To quantify utilization, we compute the average activation magnitude for each token as the mean $L_{2}$ norm of its injected vector over all occurrences and layers (Appendix C.4).
Injection demand varies across depth and modules. We next ask where token-indexed injection is actually needed. We measure, for each layer, how similar the layer’s hidden representations are to the corresponding token embeddings at the input, and we further separate the analysis by module outputs. This reveals a consistent depth-dependent drift: similarity decays with depth, indicating that token identity becomes progressively diluted as representations are repeatedly mixed. Moreover, injection demand is heterogeneous across modules, with a clear preference for modulating attention content pathways and a stronger need for injection in deeper layers. We compute similarity using cosine alignment between layer outputs and the input token embedding, averaged over tokens; the exact formula is given in Appendix C.4.
@@ B.3 Observation (extended)
Appendix C Additional Details
C.1 Injection Sites and Efficiency (extended)
@@
C.2 Additional Details of Frequency-Aware Hash Mapping
@@
C.3 Frequency-aware Routing (extended)
@@ C.5 Stability rationale
Why depth scaling uses $\sqrt{\ell+1}$ in Eq. (14). Consider a simplified residual-stream model $\mathbf{H}_{\ell}=\mathbf{H}_{\ell-1}+\mathbf{r}_{\ell}$, where $\mathbf{r}_{\ell}$ is the block update with $\mathbb{E}[\mathbf{r}_{\ell}]=\mathbf{0}$ and $\mathrm{Var}(\mathbf{r}_{\ell})\approx\sigma_{r}^{2}$ (approximately constant across depth under common initializations). Ignoring cross-layer correlations, we have $\mathrm{Var}(\mathbf{H}_{\ell})\approx(\ell+1)\sigma_{r}^{2}$, hence the RMS scale of the residual stream grows like $\sqrt{\ell+1}$[29]. Multiplying the injection gate by $\sqrt{\ell+1}$ keeps the injected residual at a comparable relative scale across depth rather than shrinking in deeper layers. The warmup factor $w_{\text{warmup}}(u)$ further prevents early optimization from being dominated by a sparsely trained injection channel.
@@
C.6 Why linear stacking does not increase capacity
@@
C.7 Why ShortConv increases expressivity

### 24 Apr 2026 v2

@@ 1 Introduction
Increasing model capacity in large language models usually comes with higher inference-time compute, making further scaling increasingly costly in practice. Token-indexed memory offers an alternative scaling path: instead of relying solely on dense computation, the model can retrieve token-conditioned features from a lookup table, thereby expanding static capacity without proportionally increasing compute [mole_ref, scone_ref]. In practice, however, prior lookup-based designs often require very large tables and substantial memory traffic, making these gains difficult to realize efficiently [mole_ref, scone_ref].
@@
To make this scaling axis practical, we ask a more basic question: what does an efficient lookup table need to optimize for under finite data and bandwidth? We analyze how token-indexed parameters are trained and how their representations behave, and find two consistent regularities. First, utilization is strongly frequency-driven under Zipf-like statistics: high-frequency items dominate updates and learned activations, while most long-tail rows remain under-trained as the table grows [yu2025scalingembeddinglayerslanguage]. Second, scaling capacity by adding parallel fixed slots is not sufficient: different slots are trained on highly overlapping events, encouraging redundant subspaces and rapid saturation [mole_ref]. These findings indicate that the bottleneck is not the lack of parameters, but the mismatch between table capacity, update density, and representational diversity.
@@ 2 General Framework
Lookup-based augmentation is fundamentally an efficiency problem: the injection design determines how much memory, activation traffic, and compute are spent per token. We therefore track three budgets: (i) memory budget (cache size) $\mathcal{C}$, dominated by the lookup table size $S\cdot d_{s}$ where $S$ is the number of physical rows and $d_{s}$ is the injected dimensionality at injection site $s$; (ii) activation budget $\mathcal{A}$, the per-token activation footprint of the injected pathway, proportional to $d_{s}$; Here, activation refers to the memory footprint of intermediate tensors saved for backpropagation, rather than nonlinear activation functions. and (iii) compute budget $\mathcal{F}$, the extra per-token FLOPs induced by lookup aggregation, feature extraction, which also scales with $d_{s}$. We refer to improvement per unit $\mathcal{C}$ as parameter efficiency, and improvement per unit $\mathcal{A}$ as activation efficiency.
@@ 3 Observation
Unless stated otherwise, we report statistics after training for 41.4B tokens; metric definitions are given in Appendix D.4.
@@
We then ask why simply adding more parallel lookup slots often yields diminishing returns even when memory budget increases. In fixed-slot designs, each slot sees almost the same token occurrences and thus receives highly correlated gradients; as a result, “more slots” does not reliably create “more directions” in representation space. We diagnose this by measuring pairwise cosine similarity between retrieved embeddings across slots (Appendix D.4). Figure 2 shows that slot representations are highly similar, indicating subspace collapse and limited angular diversity. The implication is that parameter count alone is not a sufficient scaling axis: without a mechanism that breaks symmetry across slots, additional tables mostly replicate existing features.
@@ 4.2 X-gram Extraction with Gated ShortConv
Lookup retrieval is context-independent, but whether a retrieved vector is useful depends on local context. Moreover, without a view-specific nonlinear extractor, parallel memories can collapse to redundant subspaces (Appendix D.6). We therefore refine each retrieved sequence with a lightweight, normalized, gated multi-scale ShortConv that extracts local $x$-gram cues and promotes diversity across views.
@@
Different views use different kernel sizes, so the resulting extractors collectively provide multi-scale coverage and expand the receptive field from token identity to local $x$-grams. The gate acts as a context-selective filter that suppresses collision-induced noise. View-specific kernels break symmetry across parallel memories, increasing diversity before fusion. Additional rationale is provided in Appendix D.7.
@@ 4.4 Training Stability and Optimization
Token-indexed parameters are updated sparsely and can destabilize training if their scale and learning dynamics are mismatched with the backbone. We control stability with three scale rules and a sparse-aware optimization scheme; detailed derivations are provided in Appendix D.5.
@@
For optimization, we treat lookup-table parameters as a sparse group and apply a larger effective step size than the backbone to compensate for low update frequency. Concretely, we scale the lookup learning rate as a function of expected row hit rate, which depends on table size $S$ (equivalently $\rho|\mathcal{V}|$); a practical rule is $\eta_{\text{lookup}}\propto\sqrt{S}$ up to clipping, with further refinements in Appendix D.5.
@@ 5.1 Experimental Setup
We compare against a vanilla decoder-only Transformer and representative compute-decoupled baselines that scale capacity via token-indexed parameters or embedding modules (e.g., STEM, MoLE, and SCONE-style conditional memory) [sadhukhan2026stemscalingtransformersembedding, mole_ref, scone_ref]. All methods share the same backbone at each scale (0.73B and 1.15B) and the same training token budget; differences are restricted to the token-indexed parameterization, feature extraction, and injection design. Backbone hyperparameters and training recipes are summarized in Table 9 and Appendix C.2 (extended setup in Appendix C.1).
@@
We pretrain on OLMo-mix-1124 [olmo2_paper] (web + domain data from DCLM/Dolma) [dclm_paper, dolma_paper] and evaluate with lm-evaluation-harness [eval-harness] on a fixed downstream suite with consistent prompts and shot counts (Table 12).
@@ 5.5 Related Work
A recent thread explores adding large token- or $n$-gram-indexed parameters as an auxiliary memory channel, aiming to scale capacity with minimal additional FLOPs. Mixture-of-lookup designs route tokens to a small set of retrieved vectors and combine them with lightweight mixing, enabling substantial parameter growth while keeping the runtime compute largely unchanged [mole_ref]. Similarly, scalable lookup-based conditional memory frames retrieval as a new axis of sparsity: additional capacity primarily increases memory traffic rather than matrix multiplication [scone_ref]. These methods are attractive in practice because retrieval is deterministic from token IDs and can be served from host memory with prefetching, but they also expose a core bottleneck: the efficiency of using those parameters per accessed byte, especially under Zipfian long tails where many rows are rarely updated.
@@
Another line of work enlarges the input interface by scaling embedding layers, vocabulary size, or related embedding modules [yu2025scalingembeddinglayerslanguage, huang2025overtokenizedtransformervocabularygenerally, sadhukhan2026stemscalingtransformersembedding]. Closely related work also studies memory-efficient embedding parameterizations through heterogeneous capacity allocation, compositional compression, and parameter sharing, especially in recommendation systems and large-vocabulary language models [naumov2019deep, ginart2021mixed, shi2020compositional, desai2022random, desai2022trade, desai2023defense, baevski2018adaptive, svenstrup2017hash]. These approaches mainly improve the capacity or memory efficiency of input-side embeddings, often through non-uniform allocation or shared parameterization. In contrast, we study layer-wise token-indexed memory rather than a single input embedding layer, and use frequency information to structure physical capacity allocation, with hashing applied only locally within dense buckets where collisions are unavoidable.
@@
Sparse Mixture-of-Experts (MoE) [shazeer2017outrageously] employs learnable gating to activate a sparse subset of experts per token, enabling efficient capacity scaling [lepikhin2020gshard, fedus2022switch, jiang2024mixtral]. While recent advances explore fine-grained specialization [dai2024deepseekmoe] and unified scaling laws [krajewski2024scaling], standard MoE designs typically partition experts by layer. This layer-wise isolation causes total memory to scale linearly with depth and limits the potential for cross-layer operator reuse. Existing routing strategies—ranging from load-balancing heuristics [fedus2022switch, lepikhin2020gshard] and hash-based assignments [roller2021hash] to expert-choice [zhou2022mixture], BASE layers [lewis2021base], and soft mixtures [puigcerver2024from]—primarily optimize intra-layer token allocation. However, these methods do not address the recursive, multi-layer exposure of experts, which is central to the universal expert design in X-gram.
@@
Where injection occurs matters for both stability and efficiency. Injecting into the attention value pathway can be particularly cost-effective under grouped-query attention because the value dimensionality is smaller than the query/key dimensionality, yielding lower activation and compute overhead for the same memory budget. Recent analyses of value-residual designs further highlight the importance of value-stream information flow [zhou2025valueresiduallearning]. Our method adopts value-path injection as the default site and uses depth-aware gating and normalization to keep the injected signal stable during training.
@@ 6 Conclusion
Appendix A Acknowledge
We would like to sincerely thank Jiacheng You and Songling Yang for their insightful discussions and valuable suggestions that greatly improved this work. We also extend our deep appreciation to all researchers—both publicly and anonymously involved in discussions and the review process—for their time, effort, and thoughtful feedback, which have been instrumental in refining this study.
Appendix B Notation
@@
Appendix C Full Versions of Compressed Sections
C.1 Extended Experimental Setup
We provide additional experimental details that are compressed in Section 5.1. Baselines. We compare against (i) Vanilla Transformer, the standard decoder-only backbone, and representative compute-decoupled baselines that scale capacity via embedding modules or token-indexed memory, including STEM [sadhukhan2026stemscalingtransformersembedding], MoLE [mole_ref], and SCONE [scone_ref]. All models share identical backbones (same depth/width) and are trained under the same token budget; methods differ only in the token-indexed parameterization and injection design. Data and evaluation. Models are pretrained on OLMo-mix-1124 [olmo2_paper], which integrates high-quality web text from DCLM [dclm_paper] with diverse domain-specific data from Dolma [dolma_paper]. We evaluate using lm-evaluation-harness [eval-harness] and keep prompts/shot counts fixed across methods within each benchmark (Table 12). Budget alignment. We keep additional compute from lookup and ShortConv within a small constant factor (Table 2) and align the memory budget $\mathcal{C}$ and injected views/slots across token-indexed methods (Table 11).
C.2 Training Details
@@
C.3 Observation (extended)
We conduct all observations by training two models under identical data and compute budgets: a Base Transformer and a PLE variant that injects token-indexed memory into every layer. In the PLE model, each layer aggregates $M_{\ell}$ token-indexed views and injects their normalized sum into the layer according to the PLE pipeline described in Section 4. Concretely, we use Eq. (5), and we report statistics after training for 41B tokens. We provide exact definitions for the activation norm and similarity metrics in Appendix D.4.
Trainability is dominated by Zipfian frequency. To sharpen the optimization target, we first ask a basic but overlooked question: when a token-indexed lookup table is scaled up, are its parameters actually trainable under realistic data budgets? In practice, the answer is constrained by Zipf’s law. Because token and $n$-gram sequences follow an extreme long tail, updates to large tables become highly sparse: most entries are rarely, if ever, visited, and therefore remain effectively under-trained. This under-training is not a minor edge case but a structural property of frequency distributions. For example, after processing 100M tokens, only $7.3\%$ of entries in a 2M-entry table receive more than 100 updates, whereas a 32K table reaches $97.6\%$ under the same criterion. To quantify utilization, we compute the average activation magnitude for each token as the mean $L_{2}$ norm of its injected vector over all occurrences and layers (Appendix D.4).
Injection demand varies across depth and modules. We next ask where token-indexed injection is actually needed. We measure, for each layer, how similar the layer’s hidden representations are to the corresponding token embeddings at the input, and we further separate the analysis by module outputs. This reveals a consistent depth-dependent drift: similarity decays with depth, indicating that token identity becomes progressively diluted as representations are repeatedly mixed. Moreover, injection demand is heterogeneous across modules, with a clear preference for modulating attention content pathways and a stronger need for injection in deeper layers. We compute similarity using cosine alignment between layer outputs and the input token embedding, averaged over tokens; the exact formula is given in Appendix D.4.
@@ C.3 Observation (extended)
Appendix D Additional Details
D.1 Injection Sites and Efficiency (extended)
@@
D.2 Additional Details of Frequency-Aware Hash Mapping
@@
D.3 Frequency-aware Routing (extended)
@@ D.5 Stability rationale
Why depth scaling uses $\sqrt{\ell+1}$ in Eq. (14). Consider a simplified residual-stream model $\mathbf{H}_{\ell}=\mathbf{H}_{\ell-1}+\mathbf{r}_{\ell}$, where $\mathbf{r}_{\ell}$ is the block update with $\mathbb{E}[\mathbf{r}_{\ell}]=\mathbf{0}$ and $\mathrm{Var}(\mathbf{r}_{\ell})\approx\sigma_{r}^{2}$ (approximately constant across depth under common initializations). Ignoring cross-layer correlations, we have $\mathrm{Var}(\mathbf{H}_{\ell})\approx(\ell+1)\sigma_{r}^{2}$, hence the RMS scale of the residual stream grows like $\sqrt{\ell+1}$[xiong2020layer]. Multiplying the injection gate by $\sqrt{\ell+1}$ keeps the injected residual at a comparable relative scale across depth rather than shrinking in deeper layers. The warmup factor $w_{\text{warmup}}(u)$ further prevents early optimization from being dominated by a sparsely trained injection channel.
@@
D.6 Why linear stacking does not increase capacity
@@
D.7 Why ShortConv increases expressivity